A population consists of three numbers 2, 5, 8. Enumerate all possible samples of size 2
which can be drawn without replacement from this population. Verify that the sample mean is an unbiased estimate of the population mean. Calculate the standard error of the sample
mean.
1
Expert's answer
2019-05-08T12:59:35-0400

All possible samples:{(2,5), (2,8), (5,8)}.

Sample means for these samples: 3.5, 5, 6.5.

Mean of the sample means: (3.5+5+6.5)/3=5.

Population mean: (2+5+8)/3=5.

So, the sample mean is an unbiased estimate of the population mean.

Population variance: "\sigma^2=\frac{(2-5)^2+(5-5)^2+(8-5)^2}{3}=6."

The variance of the sampling distribution of means: "\sigma_x^2=\frac{\sigma^2(N-n)}{n(N-1)}=\frac{6(3-2)}{2(3-1)}=\frac{3}{2}"

The standard deviation of the sampling distribution of means: "\sigma_x=\sqrt{\frac{3}{2}}"